StarkingYugo:

A ladder is leaning against a building, forming a 70° angle with the ground. The base of the ladder is 8.2 ft from the base of the building. What is the length of the ladder? Round your answer to the nearest tenth of a foot.

snowflake0531:

 Do you know your trigonometric ratios

snowflake0531:

e.e you're offline sin(x) = opposite/hypotenuse cos(x) = adjacent/hypotenuse tan(x) = opposite/adjacent (all from teh reference angle) So we see we have the adjacent side and the hypotenuse side, well, want to find the hypotenuse side So we use tangent tan(70) = 8.2/x So x=8.2/(tan(70))